OpenStudy (anonymous):

What are the real and complex solutions of the polynomial equation? x^4-41x^2=-400

OpenStudy (dtan5457):

x^4 - 41x^2 + 400 = 0 : standard form equation The solutions of this equation are: X1 = 4 X2 = -4 X3 = 5 X4 = -5 there are many ways to solve this, but I'd just graph it if possible and the intercepts are clearly those 4 solutions

OpenStudy (anonymous):

Can I have the steps on how you did that?

OpenStudy (dtan5457):

there are many methods but i simply typed the equation https://www.desmos.com/calculator

OpenStudy (dtan5457):

you might be able to do algebraically etc

OpenStudy (anonymous):

I figured it out but thank you!

OpenStudy (campbell_st):

this is an equation that can be written as a quadratic and then factored. \[(x^2)^2 - 41x^2 + 400 = 0\] the factors of 400 that add to -41 and -25 and -16 so in factored from its \[(x^2 -16)(x^2 - 25) = 0\] the quadratic factors are the difference of 2 squares, so can be factored and solutions found hope it helps

OpenStudy (dtan5457):

yep, that's another method^

OpenStudy (anonymous):

Technically, all the roots are complex :P

OpenStudy (campbell_st):

@uhhh how can \[x^2 - 16 = 0 ~~~and~~~~x^2 - 25 = 0\] have complex roots

OpenStudy (anonymous):

4 is a complex number and a real number -4 is a complex number and a real number 5 is a complex number and a real number -5 is a complex number and a real number

OpenStudy (anonymous):

All real numbers are complex. All imaginary numbers are complex. Sums of real and imaginary numbers are also complex.